a=b^x,b=c^y,c=a^z,then what will be thevalue of xyz ?

Grade:12

2 Answers

Nishant Vora IIT Patna
7 years ago
abc =b^x c^ya^z

hence
abc *b^-x c^-ya^-z = 1

b^1-x c^1-ya^1-z = 1 =b^0 c^0 a^0

hence this is possible only when,

x=y=z=1

Thanks
moumi roy
7 years ago
given a=b^x,b=c^y,c=a^z
then in the third criterion,xyzb^x)^z={[(c^y)^x]^z}=cc=a^z=(
so c=cxyz comparing xyz=1